OpenStudy (18jonea):

11pi/18 radians in degrees I got 110

OpenStudy (18jonea):

@confluxepic

OpenStudy (fortytherapper):

Sounds good to me atleast

OpenStudy (18jonea):

@mathstudent55

OpenStudy (fortytherapper):

Wait, since 9pi/18 is pi/2, would 10pi/18 be 110?

OpenStudy (fortytherapper):

Oh nvm, yeah 110

OpenStudy (agent0smith):

To convert from radians to degrees - multiply by 180/pi. If that's what you did, then you're correct.

OpenStudy (18jonea):

i got that but can you double check please

OpenStudy (18jonea):

@agent0smith

OpenStudy (agent0smith):

Looks fine.

OpenStudy (18jonea):

so 110? @mathstudent55

OpenStudy (18jonea):

@Hayhayz

OpenStudy (agent0smith):

Yes. You shouldn't need that much confirmation, it's a really simple calculation.

OpenStudy (mathstudent55):

\(\dfrac{11 \pi}{18} ~rad = \) \(=\dfrac{11 \pi}{18} ~rad \times \dfrac{180^\circ}{\pi ~rad} \) \(=\dfrac{11 \cancel{\pi}}{\cancel{18}} ~\cancel{rad} \times \dfrac{\cancel{180}~10^\circ}{\cancel{\pi} ~\cancel{rad}~1} \) \(= 110^\circ\)

OpenStudy (mathstudent55):

To convert from rad to deg or deg to rad use the conversion factors I gave you in your last problem.